### v1

# thanks: Corresponding author

Abstract
We introduce a method for constructing global approximations to correlation functions of strongly interacting quantum field theories, starting from perturbative results. The key idea is to employ interpolation method, such as the two-point Padé expansion, to interpolate the weak and strong coupling expansions of correlation function. We benchmark this many-body interpolation approach on two prototypical models: the lattice $\phi^{4}$ field theory and the 2D Hubbard model. For the $\phi^{4}$ theory, the resulting two-point Padé approximants exhibit uniform and global convergence to the exact correlation function. For the Hubbard model, we show that even at second order, the Padé approximant already provides reasonable characterizations of the Matsubara Green’s function for a wide range of parameters. Finally, we offer a heuristic explanation for these convergence properties based on analytic function theory.
preprint: APS/123-QED
Developing accurate and numerically efficient computational methods for strongly interactive systems remains a central challenge in quantum field theory. Non-perturbative numerical approaches, among them quantum Monte Carlo (QMC) [blankenbecler1981monte, zhang1995constrained, rubtsov2005continuous, Prokofev1998QMC, ceperley1995path, creutz1979monte, Gull11RMP], tensor-network methods [white1992DMRG, schollwock2005density, jordan2008classical], and dynamical mean-field theories [metzner1989DMFT, georges1992DMFT, georges1996dynamical, kotliar2006electronic, lichtenstein2000cluster, maier2005quantum] have provided invaluable insights into the physics of the strong correlation regime of interactive field systems, including the notoriously subtle two-dimensional Hubbard model [LeBlanc15, Arovas22, Qin22]. On the other hand, perturbative approaches [MP2, Hedin65], despite their limited accuracy in the strong correlation regime, remain attractive due to their analytic or semi-analytic nature, general applicability to electronic structure systems, and favorable scalability to large system sizes. Especially in recent years, several algorithmic developments in continuous-time quantum Monte Carlo [Gull11RMP, rubtsov2005continuous, Gull08_ctaux, Werner06, Werner06Kondo], diagrammatic Monte Carlo methods [Prokofev1998QMC, Van_Houcke_2010, rossi2017determinant, Moutenet_CDet_2018], and automatic diagram generations [kleinert2000recursive, serone2018lambdaphi4, hahn2001generating] have made it possible to evaluate the perturbation series to relatively high orders for interactive quantum fields that appear in both condensed matter and high-energy physics. With these advances, the central challenge of solving strongly interative systems with perturbative methods has gradually shifted from their computational cost to the fundamental limitations imposed by the finite radius of convergence of the perturbation series itself.
Extending the applicability of diagrammatic expansions of quantum field theories beyond the convergence radius of the original perturbation series typically requires resummation techniques. Several approaches have been proposed, including Padé approximants [baker1981pade, vsimkovic2019determinant, ferrero2023diagrammatic, serone2018lambdaphi4], resurgence theory [dorigoni2019introduction, basar2013resurgence, Rossi_resummation_2018], and action deformation strategies [Pollet_regularization_2010, Rossi_action_2016, Kim_mapping_2021]. In this paper, we propose an alternative approach based on a classical yet powerful idea from approximation theory – the two-point Padé expansion [bender2013advanced]. The general procedure is as follows: for an interacting field theory, we construct two series expansions for the correlation function, namely a weak-coupling expansion (WCE) and a strong-coupling expansion (SCE), which are valid in the limits of interaction strength $g\to 0$ and $g\to+\infty$, respectively. We then apply multi-point interpolation methods, such as the two-point Padé expansion, to interpolate between these two limits. This procedure naturally results in a global approximation to the correlation function for all $g\geq 0$.
While technically similar to the conventional one-point Padé expansion [baker1981pade, vsimkovic2019determinant, ferrero2023diagrammatic, serone2018lambdaphi4] used for WCE resummation, the proposed two-point approach differs fundamentally in two aspects. First, incorporating information from both endpoints effectively transforms the extrapolation problem into an interpolation problem, which enables the construction of an approximant that converges globally for all interaction strengths $g\geq 0$. This property is in sharp contrast to the one-point method whose convergence is inherently restricted to regimes near the origin. We numerically demonstrate this convergence in the lattice $\phi^{4}$ theory and the Hubbard model and provide a heuristic justification for the observed convergence behavior. Second, since the two-point scheme matches two limits instead of one limit, a given order of Padé approximation requires roughly half the perturbation series expansion order compared with the one-point scheme. This dramatically reduces the computational cost arising from enumerating Feynman diagrams. We note that the idea of interpolating between weak- and strong-coupling expansions has previously been applied to quantum field and string theory models in high-energy physics [sen2013s, banks2013two, pius2014s] but, to our knowledge, these applications heavily rely on duality principle to get the SCE from the WCE. A key difference here is that we employ techniques to directly obtain the strong coupling expansion of lattice field theories, which enables us to address condensed matter systems such as the Hubbard model.
A simple example.
We illustrate the central ideas using a schematic model—the zero-dimensional $\phi^{4}$ model [brown2015two]. Despite its simplicity, this example demonstrates how the two-point Padé expansion interpolates the weak and strong coupling expansion for a general field theory model and yields globally convergent approximations to the correlation function. The zero-dimensional $\phi^{4}$ model is a probability distribution over a single real variable $\phi$, with the probability density:
$\rho(\phi)\propto e^{-\frac{1}{2}m^{2}\phi^{2}-\frac{g}{4!}\phi^{4}}$ (1)
where $m,g>0$. For this interacting field, we aim to compute the two-point correlation function of the field, which is defined by its second moment: $G=\langle\phi^{2}\rangle=\frac{1}{Z}\int_{-\infty}^{+\infty}\phi^{2}\,e^{-\frac{1}{2}m^{2}\phi^{2}-\frac{g}{4!}\phi^{4}}dq$. The result admits an analytical solution:
$G=\frac{4}{m^{4}}\rho\left[\frac{K_{3/4}(\rho)}{K_{1/4}(\rho)}-1\right],$ (2)
where $\rho=\frac{3m^{4}}{4g}$ and $K_{1/4}(\cdot)$, $K_{3/4}(\cdot)$ are modified Bessel functions of the second kind. If we fix $m=1$, the function $G=G(g)$ can be approximated using two perturbative expansions. On the one hand, as $g\rightarrow 0$, the weak coupling expansion for $G(g)$ can be readily obtained by expanding $e^{-\frac{g}{4!}\phi^{4}}$ and taking the ensemble average with respect to the Gaussian weight, i.e., following the standard many-body perturbation theory (MBPT) procedure [parisi1988statistical, brown2015two]. On the other hand, the strong coupling expansion as $g\rightarrow+\infty$ can be derived in a similar manner by instead expanding the Gaussian factor $e^{-\frac{1}{2}\phi^{2}}$ and averaging with respect to the non-Gaussian weight $e^{-\frac{g}{4!}\phi^{4}}$. This leads to the following WCE and SCE series for $G(g)$:
$G(g)=$ $1-\frac{g}{2}+\frac{2g^{2}}{3}-\frac{11g^{3}}{8}+\frac{34g^{4}}{9}+\cdots$ (3)
$\begin{split}G(g)=&\left[\frac{2\sqrt{6}\,\Gamma\left(\frac{3}{4}\right)}{\Gamma\left(\frac{1}{4}\right)}\right]\frac{1}{\sqrt{g}}\\ +&12\left[\left(\frac{\Gamma\left(\frac{3}{4}\right)}{\Gamma\left(\frac{1}{4}\right)}\right)^{2}-\frac{\Gamma\left(\frac{5}{4}\right)}{\Gamma\left(\frac{1}{4}\right)}\right]\frac{1}{g}+\cdots\end{split}$ (4)
The high-order expansion coefficients can be computed numerically using recursive formulas (see the Supplemental Material (SM) [suppl]). By setting $\tilde{g}=\sqrt{g}$ and reformulating Eq. (3) and Eq. (4) into series expansions with respect to $\tilde{g}$, we can employ the two-point Padé approximant [bender2013advanced] to construct global approximations to G(g~)G($\tilde${g}). Specifically, the expansion ansatz is:
$P_{[N/N+1]}(\tilde{g})=\frac{A_{0}+A_{1}\tilde{g}+\cdots+A_{N}\tilde{g}^{N}}{1+B_{1}\tilde{g}+\cdots+B_{N+1}\tilde{g}^{N+1}},$ (5)
where the coefficients $\{A_{m},B_{m+1}\}_{m=0}^{N}$ are determined by matching the series expansions of the rational function ansatz Eq. (5) with WCE Eq. (3) and SCE Eq. (4) as:
[…]
The numerical results are summarized in Fig. 1, which demonstrates that the two-point Padé approximants converge to the exact function G(g~)G($\tilde${g}) uniformly (i.e., the error gets smaller as $N$ gets larger) and globally (i.e., it holds for all $\tilde{g}$) for the cases studied.
We note that other interpolation methods can also be used. For instance, one can directly interpolate on approximated function values obtained near the two endpoints using the truncated WCE/SCE series. Moreover, the two-point Padé ansatz also admits various generalizations. In particular, one can freely choose the matching orders of WCE and SCE and construct a general Padé-$w_{r}$-$s_{2N+1-r}$ scheme using the same rational function ansatz  Eq. (5), where $r$ and 2N+1r2N{+}1{-}r denote the matching orders of the WCE and SCE series, respectively. The two expansion points can also be taken as $\tilde{g}=0$ and $0\ll\tilde{g}=\tilde{g}_{0}<+\infty$, leading to what we shall refer to as the Padé–Taylor expansion in the remainder of this work. Detailed derivations of these variants are provided in the SM [suppl].
Figure 1: (Left) Two-point Padé approximations of G(g~)G($\tilde${g}) for the $0d$ - $\phi^{4}$ model, compared with 5-th order WCE Eq. (3) and 5-th order SCE Eq. (4) expansion results. (Right) Approximation error of two-point Padé expansions relative to the exact result Eq. (2).
Strong-coupling expansion.
[…]
The zero-dimensional $\phi^{4}$ example demonstrates the potential effectiveness of the two-point Padé scheme in constructing global approximations of correlation functions for strongly interacting fields. Once the weak- and strong-coupling expansions are available, forming the corresponding two-point Padé approximation is straightforward. In our construction, the WCE is always provided by standard MBPT [parisi1988statistical, zinn2021quantum, negele2018quantum]. Thus, the main challenge reduces to developing the SCE for interacting fields, which constitutes the most technically demanding part of our approach.
For a general interacting field, the SCE is valid in the limit of coupling strength $g\to+\infty$ and can be, in principle, obtained by treating the interaction part of the Hamiltonian as unperturbed while expanding the non-interacting part in the path integral. However, for non-zero-dimensional fields, proceeding in this naive way leads to an expansion series that quickly becomes intractable as Wick’s theorem is no longer valid for a non-quadratic unperturbed Hamiltonian. To obtain a systematic treatment of SCE, we restrict our attention to fields containing only local quartic nonlinearities, exemplified by the lattice $\phi^{4}$ field and the Hubbard model discussed below. For these systems, the Hubbard–Stratonovich transformation together with the techniques introduced by Pairault et al. [pairault1998strong, pairault2000strong] and ourselves enable us to develop a SCE framework that treats the $\phi^{4}$ field and the Hubbard model on an equal footing. This part of the derivation involves a series of combinatorial techniques which, due to their technical complexity, will be presented elsewhere. In this letter, we only make use of the final SCE results, which are summarized in the SM [suppl].
Applications.
Let us first consider the $\phi^{4}$ field theory model on a one-dimensional chain of length $N_{s}$, where the Hamiltonian is given by:
$\mathcal{H}=\sum_{ij}\frac{t_{ij}}{2}(\phi_{i}-\phi_{j})^{2}+\sum_{i}\frac{\mu}{2}\phi_{i}^{2}+\frac{g}{4!}\phi_{i}^{4}.$ (7)
[…]
Figure 3: Top-left: Padé approximation error of the Matsubara Green’s function $\frac{1}{N}\sum_{n=1}^{N}\bigl|G_{00}(i\omega_{n})-G^{\mathrm{P}}_{00}(i\omega_{n})\bigr|$ for the Hubbard dimer, using the analytic $G_{00}(i\omega_{n})$ as the reference. Other panels: Approximation results for 2D Hubbard model; solid lines show AFQMC reference and dashed lines show the Padé-$w_{2}$-$s_{1}$ approximation results. All simulations are at $t=1.0,\beta=20$. The Padé-Taylor expansion interpolates between two points: $U=0$ and $U=10$.
For the next application, we consider the Hubbard model, where the Hamiltonian is given by:
$\mathcal{H}=-\sum_{ij\sigma}t_{ij}c_{i\sigma}^{\dagger}c_{j\sigma}+\sum_{i}Un_{i\uparrow}n_{i\downarrow}-\mu\sum_{i\sigma}n_{i\sigma}$
where $t_{ij}$ denotes the nearest-neighbor hopping, and a periodic boundary condition is imposed. We restrict ourselves to the half-filling case where $\mu=U/2$ and aim to calculate the single-particle Green’s function: $G_{ij,\sigma}(\tau)=-\langle\mathcal{T}c_{i\sigma}(\tau)c_{j\sigma}^{\dagger}(0)\rangle$. The Padé interpolation is performed in frequency space, and the imaginary-time Green’s function is obtained via the Fourier transform. Again, the WCE is obtained following the standard MBPT procedure [stefanucci2013nonequilibrium, fetter2012quantum]. The SCE is obtained using procedures similar to those used for the $\phi^{4}$ theory [suppl, metzner1991linked]. However, unlike the $\phi^{4}$ case, the expansion here is carried out with respect to the hopping amplitude $|t_{ij}|$ rather than $1/U$, i.e., considering the limit $|t_{ij}|\to 0$ instead of $U\to\infty$  [^2] These two limits proves to be different for calculating the imaginary time Green’s function since the naive transformation $t\rightarrow t/U$ will accompanied with a change of inverse temperature $\beta\rightarrow\beta U$. . This introduces additional difficulty when matching the WCE and SCE series through the Padé approximation, since the rational function ansatz involves only a single variable $U$. To resolve this, a reformulation of the SCE series is introduced in the SM [suppl], which ultimately enables the construction of the Padé-Taylor expansion using $U$ as the only expansion variable for any fixed $\omega_{n}$ value.
We first benchmark our numerical results on the Hubbard dimer. The Hubbard dimer Hamiltonian is easily diagonalizable, yielding an analytic solution $G(U,i\omega_{n})$ [suppl]. From this solution, we can directly extract the WCE/SCE series via differentiation rather than using diagrammatics and construct high-order two-point Padé approximants  [^3] In this test, we temporarily set aside the diagrammatic computation problem of WCE/SCE and focus on validating the convergence property of the two-point Padé expansion. The equivalence of the WCE/SCE series obtained by differentiation and diagrammatics for the Hubbard dimer will be detailed in SM  [suppl]. As shown in the top-left panel of Fig. 3, the approximants show clear numerical convergence with increasing order. For the two-dimensional Hubbard model on a square lattice, the WCE and SCE must be evaluated diagrammatically, which we have carried out up to second order. Getting higher-order expansion series is a separate research topic. One may for example employ diagrammatic QMC methods to obtain the WCE [ferrero2023diagrammatic] and develop similar algorithms for the SCE. Nevertheless, we find that even at such low orders, the two-point Padé expansion already provides a reasonable approximation to the single-particle Green’s function over a wide range of $U/t$ values. The results are summarized in Fig. 3, where a Padé-Taylor interpolation for the Green’s function at $U=0$ and $U=10$ is performed. Specifically, for the $4\times 4$ lattice, we computed the onsite Green’s function $G_{ii}(\tau)$ and the density of states $\rho(\omega)$ via analytic continuation, and compared the results with AFQMC benchmarks [qin2016benchmark]. The Padé-$w_{2}$-$s_{1}$ approximation captures the opening of the spectral gap as well as the spread of the bands (see the bottom-left panel). We further performed calculations for larger lattices and examined $\beta G_{ii}(\tau=\beta/2)\sim A(\omega=0)$, a commonly used indicator of the metal-to-insulator transition at low temperatures. The simulation result shows that the accuracy of the approximation decreases with increasing lattice sizes, indicating that higher expansion orders are required for quantitative agreement. It awaits further investigation how the convergence rate scales as the lattice size $L\to\infty$. Since the WCE is known to exhibit insensitivity to lattice size at sufficiently high expansion orders [ferrero2023diagrammatic], one may expect a similarly favorable behavior for the SCE and hence the applicability of the two-point Padé expansion.
Figure 4: Schematic illustration of the convergence mechanism via analytic continuation. If $f(z)$ is pole-free within a $\delta$-tube enclosing $\{0\}\cup\mathbb{R}^{+}$, it can be obtained by successive AC from the germ at $z=0$ to the whole tube, thereby recovering $f(x)$ for $x\geq 0$. This extrapolation procedure may converge slowly for large $x$ (e.g. using one-point Padé expansion); however, knowing an additional germ at $0<z_{1}\leq+\infty$ allows interpolative AC between the two germs that may converge rapidly to the original function (e.g. using two-point Padé expansion). Red asterisks (*) mark the singularities of $f(z)$, and $r$ denotes the Taylor-series convergence radius at $z=0$. Note that at two ends, the series convergence regimes may be much larger and the displayed analytic germs only show their overlapping with the $\delta$-tube.
Why is the convergence possible?
Despite the complexity of obtaining the SCE for a general field-theoretic model, what makes this exploration worthwhile is the particularly striking feature that the two-point interpolation not only yields accurate approximations to the Green’s function, but also exhibits clear numerical convergence as the expansion order increases. This convergence property makes the approach especially appealing, as it offers a pathway to obtain quantitatively accurate results for strongly correlated systems in a systematically improvable and computationally scalable manner since all the calculations are based on bare perturbation results.
Proving the convergence of Padé approximants is notoriously challenging [baker1981pade, stahl1997convergence, buslaev2013convergence], let alone establishing such results for correlation functions of general statistical or quantum fields. Nevertheless, we can provide heuristic arguments to explain why such global convergence is possible. Mathematically, this amounts to asking given information about a function $f(x)$ at one or two points, whether it is possible to reconstruct the entire function for $x\in\mathbb{R}^{+}$. For a real variable $x$, this reconstruction is generally impossible due to non-uniqueness of solution. However, if $f(x)$ admits an analytic continuation to the complex plane, the situation is entirely different. As illustrated in Fig. 4, the analytically continued function $f(z)$ may possess poles or branch points in $\mathbb{C}$. But if all the singularities lie away from the real axis—say, outside a small $\delta$-tube containing $\{0\}\cup\mathbb{R}^{+}$—then $f(z)$ is uniquely defined and analytic within this domain. Consequently, starting from a locally convergent power series at $z=0$ (a germ), one can recover this globally defined analytic function through successive analytic continuation (AC)[ahlfors1979complex]. Padé expansion happens to be a practical numerical tool that can effectively perform such AC [george2012quantitative, baker1981pade]. Moreover, incorporating information from two points, i.e., employing two-point Padé expansion, effectively transforms the extrapolative AC problem into an interpolative AC problem, thereby yielding a natural global approximation to the correlation function that may converge rapidly.
The numerical convergence of the two-point Padé expansion observed for the lattice $\phi^{4}$ fields and the Hubbard model can be plausibly explained by this mechanism. In particular, for the Hubbard dimer, we can explicitly determine all poles of the analytically continued Green’s functions $G(U,i\omega_{n})$, with $U\in\mathbb{C}$   [^4] Here we particularly emphasize that the AC for the Matsubara Green’s function $G(U,i\omega_{n})$ of the Hubbard model is with respect to the interaction strength $U$ for each $i\omega_{n}$, not the AC $i\omega_{n}\rightarrow\omega+i0^{+}$ that are used to find the spectral function. , and verify that they indeed exhibit the structures shown in Fig. 4 (the analysis for $\phi^{4}$ example is similar but more delicate since the original WCE series is only asymptotic. See SM [suppl] for details). Guided by the Lee–Yang theory for the Ising model [lee1952statistical, yang1952statistical], we expect that the pole-free property near the real axis persists for finite systems at finite temperature or for infinite systems at high temperatures whenever the phase transition does not occur and the correlation functions remain analytic. Even in cases involving phase transitions, the two-point Padé expansion can still be used to probe the system’s behavior near the critical point by constructing the expansion at finite $L$ and then gradually approaching the thermodynamic limit $L\rightarrow\infty$ until the finite-size effect becomes negligible. With that being said, how precisely the strong correlation regime is approached and resolved, and how well singularities at phase transitions are approximated, is an interesting topic for future study.
Conclusion and outlook.
We developed a interpolation approach to construct global approximations to the Green’s function using the two-point Padé expansion that bridges the weak and strong coupling expansions of interacting field theories. When applied to the lattice $\phi^{4}$ field theory and the Hubbard model, the two-point Padé approximates exhibit clear convergence toward the exact correlation function across a broad range of interaction strengths. This convergence behavior can be understood through the analytic continuation of the Green’s function and its pole structure in the complex plane. As a new non-perturbative framework, this approach can be combined with CDet and other diagrammatic QMC techniques [ferrero2023diagrammatic, Van_Houcke_2010, rossi2017determinant] to obtain higher-order approximations for various correlation functions and potentially a wider class of thermodynamic quantities. We believe that this work paves a promising pathway toward accurate and efficient simulations of strongly correlated systems, potentially offering new insights into long-standing problems in the relevant field.
We thank S. Zhang’s group for providing us the AFQMC results for the 2D Hubbard model. Y. Zhu acknowledges valuable discussions with colleagues, including L. Lin, S. Zhang, Z. Huang, P. Xie, P. Rosenberg and V. Vlcek. Y. Yu acknowledges a helpful discussion with Jiuci Xu. This work is supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research and Office of Basic Energy Sciences, Scientific Discovery through Advanced Computing (SciDAC) program under Award Number DE-SC0022198. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award BES-ERCAP0029462 (project m4022) and ASCR-ERCAP-m1027. Y.Yu was supported by the National Science Foundation under Grant No. NSF DMR 2401159. EK was supported by the U.S. Department of Energy (DOE) under Contract No. DE-AC02-05CH11231 through the Office of Advanced Scientific Computing Research Accelerated Research for Quantum Computing Program, FAR-QC. EG acknowledges funding from the European Research Council (ERC) under Advanced Grant No. 101142136 (Quantum Algorithms).